12. Write sin (x - Tt) in simplified form.
Answer #1

Sin (2-1) = sin (-(A-2) for --Sin (~-~9 ( Jin (0) = -tino) -Sina Jin (a-n) (ra = Sinn nAlternative we have sin (A-B)=sin A co/ B - LosAsin B Jinlar M =fina cosa Sind .(-1) Cos (29.0 cosa=1, - Ffinx (rean cosa sin